OpenStudy (anonymous):

Find the area of a regular pentagon with side length = 12 and an apothem = 3.

OpenStudy (anonymous):

45 90 108 360

jhonyy9 (jhonyy9):

90 i think because will get this from 12*3/2 =18*5=90

OpenStudy (anonymous):

That's what I was thinking, too. So I think I'm going to go with 90

jhonyy9 (jhonyy9):

ok good luck bye